< Question 1 of 19 > What concentration of Soz- is in equilibrium with Ag, Soz(s) and 9.30 x 10-3 M Ag+? The Ksp of Ag, Soz cKsp of Ag2SO3 is 1.50×10–14
Answer #1

Ag2SO3 <-----------> 2Ag+ + SO3-2

Ksp = [Ag+]2 [SO3-2]

Ksp = 1.50 x 10-14

[Ag+] = 9.30 x 10-3 M

[SO3-2] = ?

1.50 x 10-14 = [9.30 x 10-3]2 [SO3-2]

1.50 x 10-14 = 0.0000865 [SO3-2]

[SO3-2] = 1.73 x 10-10 M